### v7

# DreamBarbie: Text to Barbie-Style 3D Avatars

Abstract
To integrate digital humans into everyday life, there is a strong demand for generating high-quality, fine-grained disentangled 3D avatars that support expressive animation and simulation capabilities, ideally from low-cost textual inputs. Although text-driven 3D avatar generation has made significant progress by leveraging 2D generative priors, existing methods still struggle to fulfill all these requirements simultaneously. To address this challenge, we propose DreamBarbie, a novel text-driven framework for generating animatable 3D avatars with separable shoes, accessories, and simulation-ready garments, truly capturing the iconic “Barbie doll” aesthetic. The core of our framework lies in an expressive 3D representation combined with appropriate modeling constraints. Unlike prior methods, we use G-Shell to uniformly model watertight components (e.g., bodies, shoes) and non-watertight garments. By reformulating boundaries as Euclidean field intersections instead of manifold geodesics, we propose an SDF-based initialization and a hole regularization loss that together achieve a $100\times$ speedup and stable open topology without image input. These disentangled 3D representations are then optimized by specialized expert diffusion models tailored to each domain, ensuring high-fidelity outputs. To mitigate geometric artifacts and texture conflicts when combining different expert models, we further propose several effective geometric losses and strategies. Extensive experiments demonstrate that DreamBarbie outperforms existing methods in both dressed human and outfit generation. Our framework further enables diverse applications, including apparel combination, editing, expressive animation, and physical simulation. Project page: https://xiaokunsun.github.io/DreamBarbie.github.io/.
I Introduction
Automating the creation of 3D avatars without manual effort remains a critical challenge in computer vision and graphics, with applications ranging from VR/AR and virtual try-ons to gaming and beyond [53, 15, 70]. To fully unlock the potential of these applications, it is essential that the generated avatars meet several key standards: (1) High Quality: Digital avatars should feature exquisite geometry and realistic appearances; (2) Fine-Grained Decoupling: Enabling separation of body, clothing, shoes, and accessories within the avatar model for flexible customization and editing; (3) Expressive Animation: Supporting animations through body movements, facial expressions, and hand gestures; (4) Simulation Compatibility: Ensuring that digital avatars dress in non-watertight garments, integrating seamlessly with existing physical simulation pipelines. In vivid terms, we expect the generated digital human with qualities reminiscent of Barbie dolls, thereby enhancing their practical value across application scenarios.
Recently, text-driven 3D avatar generation methods [19, 29, 22], which integrate pretrained text-to-image (T2I) diffusion models [56, 58], have gained significant attention by eliminating the need for human image/video/3D data collection [59, 1, 36, 57, 71, 8]. However, as summarized in Table I, existing text-to-avatar methods fail to meet the above “Barbie-like” standard. Specifically, approaches [19, 29, 23, 78, 3, 21, 68, 12] leveraging implicit NeRF [46, 67] cannot support physical simulation or accurately model facial expressions and hand movements due to a lack of explicit structures. Moreover, these methods tend to produce overly smooth surfaces. Methods [41, 76, 24, 16, 37, 39, 45, 4, 66] using explicit 3DGS [27] and SMPL-X [51] enhance animation expressiveness and simulation compatibility but struggle to capture delicate geometric details. Approaches [22, 73] based on the hybrid representation DMTet [61], integrating strengths of both implicit and explicit 3D representations (i.e., stable optimization properties and explicit structure), are capable of modeling detailed geometry. However, DMTet cannot generate non-watertight surfaces required for realistic garment simulation. Furthermore, current text-to-disentangled-avatar methods typically adopt a single general diffusion model to guide both body and apparel generation. This general constraint compromises in-domain fidelity and fails to produce diverse outfits (e.g., shoes, necklaces, glasses, or other accessories).
It is natural to ask: Can we automatically generate Barbie-like 3D avatars based on text? Our answer is yes, but this is a non-trivial task due to several challenges. (1) How can we model both watertight components (e.g., bodies, shoes, and accessories) and non-watertight, simulation-ready garments within a unified framework? While more expressive representations (e.g., G-Shell [42]) offer potential solutions, initializing and regularizing them to accurately represent open surfaces without multi-view image inputs remains an open problem. (2) Providing suitable constraints for optimizing decoupled bodies and outfits to achieve domain-specific realism is both key and challenging.
TABLE I: Comparisons of Existing Text-Driven 3D Avatar Generation Methods
Methods Representation High Quality Fine-Grained Decoupling Expressive Animation Simulation Compatibility
[19, 29, 23, 78, 3] NeRF/Neus
[21, 68, 12] NeRF/Neus
[41, 76] 3DGS
[24] 3DGS
[16] 3DGS
[37] 3DGS
[39, 45, 4] SMPL-X
[66] SMPL-X
[22, 73, 80] DMTet
DreamBarbie (Ours) G-Shell
  • $\dagger$ Although SimAvatar [37] does not separate the body, clothing, shoes, and accessories, it disentangles the body, clothing, and hair. Therefore, we consider it to have achieved fine-grained decoupling.
To address these challenges, we introduce DreamBarbie [^1] Barbie® is a registered trademark of Mattel, Inc. This project is for academic research purposes only and is not affiliated with, endorsed by, or connected to Mattel, Inc., a novel text-driven framework for generating Barbie-style 3D avatars. As shown in Fig. LABEL:fig:teaser, avatars from DreamBarbie not only exhibit exquisite geometry and textures but also feature a variety of decoupled shoes, accessories, and simulation-ready garments that can be freely combined and edited. Furthermore, our framework supports expressive animation and compatibility with physical simulations, thereby satisfying all four aforementioned requirements simultaneously.
The core of the framework consists of two key components: (1) Unlike prior methods, we utilize expressive G-Shell [42] to unify watertight and non-watertight modeling. We propose an SDF-based boundary initialization that avoids manifold geodesics via a 3D proxy (Pie Mesh), ensuring stable gradients and $100\times$ acceleration. Combined with a hole-preserving loss, this enables clean open surface modeling without multi-view supervision. (2) Instead of relying on a single general model as in prior text-to-disentangled-avatar methods, we appropriately incorporate different expert diffusion models to guarantee the domain-specific fidelity. A series of effective regularization losses and strategies is also proposed to address geometric artifacts and texture conflicts when combining different expert models. Specifically, our method generates Barbie-like 3D digital humans in three stages: First, we generate a reasonable and realistic base body using human-specific diffusion models along with a proposed SMPLX-evolving prior loss. Second, we initialize apparel with the semantic-aligned body and optimize it using object-specific generative priors and several solid geometric losses. Finally, we jointly fine-tune the assembled avatar to enhance texture harmony and consistency. Through this pipeline, the generated 3D avatars are animatable and disentangled, featuring high-quality bodies, shoes, accessories, and simulation-ready garments, delivering a truly immersive “Barbie-style” digital experience.
Our main contributions are summarized as follows:
  • We introduce DreamBarbie, a novel text-driven framework for generating realistic and highly disentangled 3D avatars. The framework enables the decoupling of bodies, garments, shoes, and accessories, while enabling outfit transfer, editing, expressive animation, and simulation.
  • We innovatively adopt G-Shell to uniformly model watertight and non-watertight components with high fidelity. We further propose an efficient initialization strategy and a hole-preserving loss to ensure clean open surface modeling. To the best of our knowledge, this is the first work to incorporate G-Shell into text-to-3D generation.
  • We effectively integrate expert models at different optimization stages to provide suitable guidance, significantly improving the in-domain realism of generated components. Additionally, we introduce a series of geometric losses and strategies to address geometric artifacts and texture conflicts when combining different expert models.
  • Extensive experiments demonstrate that DreamBarbie outperforms existing methods in avatar and outfit generation, achieving superior results in geometry quality, texture detail, and alignment with text descriptions.
II Related Work
Text to Holistic 3D Avatar Generation. The field of text-to-3D generation has witnessed significant progress [52, 7, 75, 65, 69, 38, 54, 62], largely driven by advances in text-to-image models [56, 58, 79]. Building upon the success in generating general 3D objects, various methods [19, 29, 23, 78, 3, 41, 76, 24, 39, 45, 4, 22, 73] have been proposed to generate complex 3D avatars. AvatarCLIP [19] pioneers zero-shot generation of 3D digital humans from text prompts by leveraging human priors and CLIP [55]. With the introduction of the Score Distillation Sampling (SDS) loss [52], several subsequent works [29, 23, 78, 3] combine it with parametric human models [43, 51, 2] to significantly improve the fidelity of generated avatars. To enhance animation expressiveness and simulation compatibility, approaches [41, 76, 24, 39, 45, 4] replace implicit NeRF [46, 67] with explicit representations such as 3DGS [27] and SMPL-X [51] for modeling 3D humans. However, purely implicit or explicit representations struggle to capture complex geometric details, limiting the quality of generated digital humans. Accordingly, HumanNorm [22], SeeAvatar [73], and AvatarStudio [80] propose using the hybrid representation DMTet [61], which combines the advantages of both implicit and explicit representations to achieve high-fidelity avatar generation. Nevertheless, these methods typically model the human body, garments, shoes, and accessories as a single model, sacrificing flexibility and controllability for downstream applications such as apparel composition and physical simulation.
Text to Disentangled 3D Avatar Generation. To enable controllable avatar creation, several disentangled approaches have been proposed that separately model the human body and outfits through multi-stage optimization. Humancoser [68] and AvatarFusion [21] use two implicit NeRF [46, 67] to represent the body and clothing, respectively. LAGA [16] and TELA [12] adopt a layer-wise framework, modeling each clothing item as an independent layer, further improving generation flexibility and controllability. SO-SMPL [66] and SimAvatar [37] leverage sequentially offset SMPL-X [51] and compositional 3DGS [27] anchored on the mesh to enhance compatibility with simulation pipelines. However, these methods rely on a single general T2I model to guide both body and clothing generation, leading to compromised fidelity in geometry or texture within specific domains. Moreover, they struggle to generate diverse accessories such as shoes, necklaces, glasses, or other fine-grained items.
In contrast, avatars from DreamBarbie not only exhibit exquisite geometry and appearance but also wear multiple separable, realistic outfits, while supporting expressive animation and physical simulation. We summarize the main differences between DreamBarbie and existing methods in Table I.
3D Representation of Decoupled Avatar Modeling. The disentangled modeling of the human body and clothing has been extensively studied in computer vision and graphics. Depending on specific problem settings, different representations are chosen based on trade-offs among optimization properties, geometric modeling capabilities, animation expressiveness, and simulation compatibility: (1) Implicit representations [46, 67, 49, 10] are widely adopted in decoupled 3D avatar reconstruction and generation tasks [68, 21, 12, 6] due to their stable optimization properties. However, they lack explicit structures, making them unsuitable for expressive animation and physical simulation. (2) Explicit representations [27, 43, 51] are commonly used in disentangled digital human creation [16, 37, 66, 26, 40], as they provide explicit structures suitable for animation and simulation. Nevertheless, they struggle to capture delicate geometric details. (3) Combinations of the above representations have also been proposed to achieve high-fidelity reconstruction and generation of decoupled 3D avatars [14, 77, 13, 11]. Despite these efforts, the inherent disadvantages mentioned above persist. (4) Hybrid representations [61, 42] combine the advantages of implicit and explicit representations to sculpt impressive geometric details in reconstructed disentangled 3D avatars [28, 5]. G-Shell [42], in particular, addresses the limitation that DMTet [61] cannot model non-watertight surfaces compatible with physical simulations. However, how to initialize and regularize G-Shell without strong multi-view image input remains an open problem. In this work, we aim to unlock the potential of G-Shell for the challenging task of text-driven Barbie-style 3D avatar generation.
III Method
III-A Preliminary
SMPL-X [51] is a parametric human model that represents the shape, pose, and expression using a set of parameters. Given shape parameters $\beta$, body pose parameters ${\theta}_{body}$, jaw pose parameters ${\theta}_{jaw}$, hand pose parameters ${\theta}_{hand}$, and expression parameters $\psi_{exp}$, it generates a 3D human mesh ${M}_{smplx}$.
Score Distillation Sampling [52] leverages a pre-trained T2I model to guide the alignment of a 3D representation with input text. Given a text prompt $y$, a 3D representation parameterized by $\theta$, and a diffusion model parameterized by $\phi$, the SDS loss is defined as:
$\nabla_{\theta}\mathcal{L}_{SDS}=\mathbb{E}_{t,\epsilon}\left[w(t)({\epsilon}_{\phi}(x_{t};y,t)-\epsilon)\frac{\partial x}{\partial\theta}\right],$ (1)
where $t$ is the time step in the 2D diffusion process, $x=g(\theta)$ is the image rendered from $\theta$ by a differentiable renderer [30] $g(\cdot)$, $x_{t}=x+\epsilon$ is a noised version of $x$, ${\epsilon}_{\phi}(x_{t};y,t)$ is the denoised image, and $w(t)$ is the weight function. For simplicity, we omit $w(t)$ in the following formulas.
DMTet [61] is a hybrid representation that combines an implicit signed distance field (SDF) $s:\mathbb{R}^{3}\mapsto\mathbb{R}$ and a differentiable Marching Tetrahedral layer ${DMT}(\cdot)$ to transfer an implicit SDF into an explicit watertight mesh $M_{wt}$. This process is formulated as: $M_{wt}={DMT}(s(q))$, where $q$ is a set of predefined sampled points. By integrating the strengths of both implicit and explicit 3D representations (i.e., stable optimization properties and explicit structure), DMTet enables high-fidelity geometric modeling.
G-Shell [42] is an expressive representation capable of modeling both watertight and non-watertight surfaces. Building upon DMTet, G-Shell introduces a manifold signed distance field (mSDF) $\hat{s}:\mathbb{R}^{3}\mapsto\mathbb{R}$ on a watertight template. In this field, the sign indicates whether a point lies on an open surface, while the absolute value represents the geodesic distance to the boundary. A non-watertight mesh $M_{nwt}$ can then be extracted through the G-Shell mesh extraction process ${GSE}(\cdot)$: $M_{nwt}={GSE}(s(q),\hat{s}(q))$. Additionally, G-Shell can also model a watertight mesh $M_{wt}$ by ignoring mSDF. In this case, G-Shell is equivalent to DMTet: $M_{wt}={GSE}(s(q))$.
III-B Overview
Given a text prompt, DreamBarbie aims to create an animatable, disentangled 3D avatar dressed in simulation-ready garments, along with diverse shoes and accessories, resembling iconic Barbie dolls. Specifically, our framework consists of three stages: (1) Human Body Generation: This stage generates a reasonable and realistic basic human body by leveraging human-specific generative priors and a novel SMPLX-evolving prior loss (Sec. III-C and Fig. 2); (2) Apparel Generation: This stage models high-quality garments, shoes, and accessories piece by piece, utilizing object-specific diffusion models together with several initialization strategies and geometric losses (Sec. III-D and Fig. 3); and (3) Unified Texture Refinement: This stage enhances visual harmony and consistency by jointly fine-tuning the composed avatar (Sec. III-E and Fig. 3).
Figure 2: The process for generating a basic human model involves two steps: (a) Employing human-specific geometry-aware diffusion models and the SMPLX-evolving prior loss to model realistic and reasonable body shapes. (b) Subsequently, using a normal-conditioned diffusion model to generate lifelike human textures.
III-C Human Body Generation
Since the human body is a closed surface, we utilize G-Shell ${\theta}_{h}$ in watertight mesh mode (i.e., ignoring mSDF and retaining only SDF) and a texture field ${\psi}_{h}$ to model the geometry and appearance of the human body, respectively. These components are optimized under human-specific generative priors and a novel SMPLX-evolving prior loss, as illustrated in Fig. 2.
Human Body Geometry Modeling. To enhance the robustness of 3D human generation, we first optimize the shape parameters $\beta$ of an SMPL-X model using SDS loss, aligning the coarse body shape with the input text description. The resulting initial mesh, denoted $M_{init}$, serves two key purposes: (1) initializing the human G-Shell $\theta_{h}$, and (2) providing rich, semantically aligned human priors to guide realistic outfit generation. As discussed in the introduction, using a general T2I model struggles to provide domain-specific constraints necessary for creating realistic human bodies or outfits. To address this limitation, we employ human-specific diffusion models from HumanNorm [22], which are fine-tuned with high-fidelity human data.
In particular, the human-specific diffusion models include a normal-adapted diffusion model ${\phi}_{hn}$, a depth-adapted diffusion model ${\phi}_{hd}$ for human shape generation, and a normal-conditioned diffusion model ${\phi}_{hc}$ for human texture creation. The geometry-aware diffusion models optimize the initialized human G-Shell ${\theta}_{h}$ using the following SDS losses:
$\nabla_{{\theta}_{h}}\mathcal{L}^{{hn}}_{SDS}$ $=\mathbb{E}_{t,\epsilon}\left[({\epsilon}_{{\phi}_{hn}}(n^{h}_{t};y_{h},t)-\epsilon)\frac{\partial n^{h}}{\partial{\theta}_{h}}\right],$ (2)
$\nabla_{{\theta}_{h}}\mathcal{L}^{{hd}}_{SDS}$ $=\mathbb{E}_{t,\epsilon}\left[({\epsilon}_{{\phi}_{hd}}(d^{h}_{t};y_{h},t)-\epsilon)\frac{\partial d^{h}}{\partial{\theta}_{h}}\right],$ (3)
where $n^{h}$ and $d^{h}$ are the rendered normal and depth maps of the human body, respectively. $y_{h}$ denotes the input minimal-clothed human body description.
SMPLX-Evolving Prior Loss. Although Eq. 2 and Eq. 3 enable the creation of intricate human bodies, the overly strong human-specific generative priors cause the generated results to overfit the input text. This leads to unnatural geometry and exaggerated body proportions (Fig. 8-(a)). A straightforward approach is to introduce parametric human body models (e.g., SMPL-X) and provide human body priors using the following equation:
$\mathcal{L}_{{prior}}=\sum_{p\in P}\left\|s_{\theta_{h}}(p)-s_{{init}}(p)\right\|_{2}^{2},$ (4)
where $s_{\theta_{h}}(\cdot)$ and $s_{{init}}(\cdot)$ represent the SDF of the generated body G-Shell ${\theta}_{h}$ and the initialized SMPL-X mesh $M_{init}$, respectively, and $P$ is a set of randomly sampled points in space. However, due to the overly smooth geometry of SMPL-X, this approach may lead to a reduction in diversity and fine details (Fig. 8-(a)).
To address this problem, we propose the SMPLX-evolving prior loss inspired by the evolving constraint introduced in SeeAvatar [73]. Specifically, we freeze all parameters of the SMPL-X model and enhance the initial mesh $M_{init}$ into an evolvable representation $\hat{M}_{init}$ by adding learnable vertex-wise offsets. These offsets are periodically fitted to the current body every $\delta$ iteration, allowing the optimized offsets to model complex body features. The fitting objective is as follows:
$\mathcal{L}_{fit}=\lambda_{{chamf}}\mathcal{L}_{chamf}+\lambda_{{edge}}\mathcal{L}_{edge}+\lambda_{{nor}}\mathcal{L}_{nor}+\lambda_{{lap}}\mathcal{L}_{lap},$ (5)
where $\mathcal{L}_{chamf}$ is the Chamfer distance between $\hat{M}_{init}$ and the current body shape, $\mathcal{L}_{edge}$, $\mathcal{L}_{nor}$ and $\mathcal{L}_{lap}$ are the edge length regularization, normal consistency loss, and Laplacian smoothness, respectively. The advantages of this loss are twofold: (1) Enhanced Modeling Capabilities: Compared to the smooth $M_{init}$, our $\hat{M}_{init}$ excels at capturing detailed geometric features, such as muscle contours (see the zoomed area at the bottom of Fig. 2), and offers reliable yet diverse priors for subsequent geometry generation (Fig. 8-(a)). (2) Topology Preservation: Unlike the evolving templates with arbitrary topology in SeeAvatar [73], our $\hat{M}_{init}$ preserves the SMPL-X model’s topology and semantics during the evolution process. This makes it particularly suitable for apparel initialization, composition, animation, and simulation (Fig. 9).
In summary, the loss function for optimizing the human body geometry is formulated as follows:
$\mathcal{L}_{{hum-geo}}=\mathcal{L}^{{hn}}_{SDS}+\mathcal{L}^{{hd}}_{SDS}+\lambda_{{prior}}\mathcal{L}_{{prior}}.$ (6)
Figure 3: The process for generating apparel involves three steps: (a) Initializing apparel with the semantic-aligned human body. (b) Modeling apparel piece by piece using object-specific generative priors and geometric losses. (c) Refining the texture of the assembled avatar using a unified texture refinement process.
Human Body Texture Modeling. Given the human mesh generated from the previous stage, we fix it and utilize a texture field ${\psi}_{h}$ which maps a query position to its color to generate a normal-aligned body appearance. This field is optimized using a normal-conditioned diffusion model ${\phi}_{hc}$ with loss defined as:
$\nabla_{{\psi}_{h}}\mathcal{L}^{{hc}}_{SDS}=\mathbb{E}_{t,\epsilon}\left[({\epsilon}_{{\phi}_{hc}}(c^{h}_{t};n^{h},y_{h},t)-\epsilon)\frac{\partial c^{h}}{\partial{\psi}_{h}}\right],$ (7)
where $c^{h}$ represents the rendered color image of the generated human body. Since the vanilla SDS loss often leads to color oversaturation, we replace it with the following multi-step SDS (MSDS) loss [22] to further enhance the texture’s realism in later iterations of texture optimization:
$\nabla_{{\psi}_{h}}\mathcal{L}^{{hc}}_{MSDS}=\mathbb{E}_{t,\epsilon}\left[(h(c^{h}_{t};n^{h},y_{h},t)-\epsilon)\frac{\partial c^{h}}{\partial{\psi}_{h}}\right]+$ (8)
$\mathbb{E}_{t,\epsilon}\left[(V(h(c^{h}_{t};n^{h},y_{h},t))-V(\epsilon))\frac{\partial V(\epsilon)}{\partial\epsilon}\frac{\partial c^{h}}{\partial{\psi}_{h}}\right],$
where $V$ denotes the first $k$ layers of the VGG network [64], and $h(\cdot)$ represents the multi-step image generation function of the normal-aligned diffusion model.
III-D Apparel Generation
Given an SMPLX-aligned basic body generated in the previous stage, we proceed to dress the Barbie-like avatar in various outfits. Similar to body modeling, we employ G-Shell ${\theta}_{a}$ for apparel geometry and a texture field ${\psi}_{a}$ for apparel texture. Closed shoes and accessories are represented using G-Shell in watertight mesh mode (i.e., ignoring mSDF and retaining only SDF), while open garments are modeled using G-Shell in non-watertight mesh mode (i.e., retaining both SDF and mSDF). As illustrated in Fig. 3, the apparel generation process begins by initializing apparel using predefined SMPL-X masks that cover over a dozen daily outfits. We then create high-quality apparel piece by piece using object-specific generative priors and well-designed geometric losses. Lastly, we fine-tune the assembled avatar to improve appearance harmony and consistency.
Figure 4: (a) Closed Surface Initialization: Expand and sew the open mesh (cropped via SMPL-X mask) to create a closed template mesh $M_{temp}$, used to initialize the SDF $s_{\theta_{a}}(\cdot)$ of ${\theta}_{a}$. (b) Open Surface Initialization: Fit a watertight pie mesh $M_{pie}$ over the holes of the cropped open mesh. Use its SDF values to initialize the mSDF $\hat{s}_{\theta_{a}}(\cdot)$ of ${\theta}_{a}$. (c) Comparison: Contrast geodesic-based and SDF-based mSDF initialization.
Apparel Initialization. In the absence of multi-view supervision [42, 5], text-driven optimization of G-Shell representations heavily depends on effective initialization, especially for open surface modeling. Therefore, we carefully design two strategies for initializing open clothes and closed shoes/accessories, enabling accurate topology modeling for each category. (1) Closed Surface Initialization: We crop the open sub-mesh of the human body using the corresponding SMPL-X mask, expand it along vertex normals to create inner and outer layers, and sew them to form a double-layer closed template mesh $M_{temp}$. This mesh initializes the SDF $s_{\theta_{a}}(\cdot)$ of the apparel G-Shell ${\theta}_{a}$. Fig. 4-(a) shows this process. (2) Open Surface Initialization: As mSDF requires a closed manifold domain, we first construct a single-layer watertight template, $M_{temp}$, by hole-filling the cropped body mesh. This simplified topology serves as the canonical domain for the mSDF. The primary challenge lies in defining the open boundaries on $M_{temp}$. While a geodesic-based initialization (i.e., signed geodesic distance to boundaries) is intuitive, it is computationally prohibitive ($\sim$5 hours) and mathematically ill-posed for 3D implicit fields. Because geodesics are confined to the 2-manifold, they lack a stable gradient flow in 3D space, resulting in optimization artifacts (Fig. 4-(c)).
To address this, we reformulate boundary modeling as a field-based intersection problem. Specifically, we introduce the Pie Mesh ($M_{pie}$), a localized watertight proxy that encapsulates each hole’s boundary (see Supp. Mat. for fitting details). Mathematically, instead of manifold traversal, we use $M_{pie}$’s SDF values to initialize the mSDF $\hat{s}_{\theta_{a}}(\cdot)$. This ensures that points sampled inside $M_{pie}$ have negative mSDF values (indicating they lie inside the open surface), while those outside have positive values (indicating they lie outside the open surface). Consequently, the open surface is implicitly determined by the set subtraction between the $M_{temp}$ manifold and the $M_{pie}$ volume (Fig. 4-(b), black dotted box). Fig. 4-(b) illustrates the whole process. This strategy offers three advantages: (1) Efficiency: Replacing manifold-constrained path-finding with SDF queries achieves a $100\times$ speedup ($\sim$3 mins vs. $\sim$5 hours, see the Supp. Mat. for speed comparison details). (2) Stability: The SDF provides a globally continuous gradient field, ensuring stable supervision and eliminating floating artifacts or unsmooth boundaries. (3) Regularization: $M_{pie}$ serves as a geometric prior, providing explicit boundary constraints for subsequent optimization (see Eq. 14). Fig. 4-(c) compares these two initialization methods.
Apparel Geometry Modeling. Similar to human body geometry generation, we utilize human-specific diffusion models to optimize outfit geometry with the SDS loss:
$\nabla_{{\theta}_{a}}\mathcal{L}^{{hn}}_{SDS}$ $=\mathbb{E}_{t,\epsilon}\left[({\epsilon}_{{\phi}_{hn}}(n^{a+h}_{t};y_{a+h},t)-\epsilon)\frac{\partial n^{a+h}}{\partial{\theta}_{a}}\right],$ (9)
$\nabla_{{\theta}_{a}}\mathcal{L}^{{hd}}_{SDS}$ $=\mathbb{E}_{t,\epsilon}\left[({\epsilon}_{{\phi}_{hd}}(d^{a+h}_{t};y_{a+h},t)-\epsilon)\frac{\partial d^{a+h}}{\partial{\theta}_{a}}\right],$ (10)
where $y_{a+h}$ is the text description of the clothed avatar, and $n^{a+h}$ and $d^{a+h}$ are the rendered clothed human normal and depth maps, respectively. However, human-specific diffusion models are not well-suited for creating high-quality clothes, shoes, and accessories because they are not good at modeling general 3D objects. Hence, we additionally introduce the object-specific diffusion models [54] pretrained on the LAION dataset [60] for providing in-domain details and diversity.
Concretely, these diffusion models include a normal-depth diffusion model ${\phi}_{ond}$ for optimizing the apparel geometry, and a depth-conditional diffusion model ${\phi}_{oc}$ for creating the apparel texture. The normal-depth model provides effective supervision for outfit geometry generation by accurately modeling the joint distribution of normal and depth maps using the following SDS loss:
$\nabla_{{\theta}_{a}}\mathcal{L}^{{ond}}_{SDS}$ $=\mathbb{E}_{t,\epsilon}\left[({\epsilon}_{{\phi}_{ond}}(n^{a}_{t};y_{a},t)-\epsilon)\frac{\partial n^{a}}{\partial{\theta}_{a}}\right]$ (11)
$+\mathbb{E}_{t,\epsilon}\left[({\epsilon}_{{\phi}_{ond}}(d^{a}_{t};y_{a},t)-\epsilon)\frac{\partial d^{a}}{\partial{\theta}_{a}}\right],$
$\nabla_{{\theta}_{a}}\mathcal{L}^{{sdn}}_{SDS}=\mathbb{E}_{t,\epsilon}\left[({\epsilon}_{{\phi}_{sd}}(n^{a}_{t};y_{a},t)-\epsilon)\frac{\partial n^{a}}{\partial{\theta}_{a}}\right],$ (12)
where $y_{a}$ is the description of a single outfit, and $n^{a}$ and $d^{a}$ are the rendered normal and depth maps of the apparel in the undressed state. Besides, $\mathcal{L}^{{sdn}}_{SDS}$ is the vanilla Stable Diffusion (SD) SDS loss enforced on the rendered apparel normal maps. As shown in RichDreamer [54], naive SD helps object-specific diffusion models produce more stable results. In this way, two specific generative priors provide powerful and reasonable guidance to modeling high-fidelity outfits (Fig. 8-(b)).
Figure 5: Diverse Range of Barbie-Style Avatar Generation. Rendering color images and normal images for visualization. Please zoom in to see the details and see Supp. Mat. for video results.
Template-Preserving Loss. Similar to human-specific diffusion models, relying on object-specific diffusion models may lead to geometric artifacts (e.g., unexpected holes, see Fig. 8-(d)). To address this problem, we propose a template-preserving loss that enforces the apparel geometry to fit the template mesh $M_{temp}$, ensuring geometric integrity. This loss is formulated as follows:
$\mathcal{L}_{{temp}}=\sum_{p\in P}\left\|s_{\theta_{a}}(p)-s_{{temp}}(p)\right\|_{2}^{2},$ (13)
where $s_{temp}(\cdot)$ represents the SDF of the template mesh $M_{temp}$, and $P$ is the set of randomly sampled points in space.
Hole-Preserving Loss. Since the object-specific diffusion model is primarily trained on watertight mesh data, it tends to guide the G-Shell to model closed surfaces, resulting in many floating triangles inside the holes (Fig. 8-(e)). Moreover, due to the absence of multi-image input, we cannot reliably supervise the open surface like other G-Shell reconstruction works [42, 5]. To address this issue, we propose a hole-preserving loss that exploits the pie mesh $M_{pie}$ obtained from the SDF-based mSDF initialization. This loss is formulated as:
$\mathcal{L}_{{hole}}=\sum_{p\in P}\left\|\hat{s}_{\theta_{a}}(p)-s_{{pie}}(p)\right\|_{2}^{2},$ (14)
where $s_{{pie}}(\cdot)$ represents the SDF of the pie mesh $M_{pie}$. This loss effectively suppresses floating triangles by using the SDF of $M_{pie}$ to supervise the mSDF of ${\theta}_{a}$, thereby preserving clean and well-defined hole structures.
Collision Loss. To ensure the generated apparel does not intersect with the underlying human body (Fig. 8-(f)), we introduce the following collision loss $\mathcal{L}_{{colli}}$:
$\mathcal{L}_{{colli}}=\sum_{v\in V}\text{relu}(-s_{h}(v)),$ (15)
where $V$ is the set of vertices of the generated apparel mesh, and $s_{h}(\cdot)$ is the SDF of the human body.
In summary, the loss function for optimizing the apparel geometry is as follows:
$\mathcal{L}_{{app-geo}}=\mathcal{L}^{{hn}}_{SDS}+\mathcal{L}^{{hd}}_{SDS}+\mathcal{L}^{{ond}}_{SDS}+\mathcal{L}^{{sdn}}_{SDS}$ (16)
$+\lambda_{{temp}}\mathcal{L}_{{temp}}+\lambda_{{hole}}\mathcal{L}_{{hole}}+\lambda_{{colli}}\mathcal{L}_{{colli}},$
where $\mathcal{L}_{{hole}}$ is only used in non-watertight mesh mode.
Apparel Texture Modeling. Given the generated apparel shape and the textured human body, we employ an object-specific depth-conditional diffusion model for lifelike apparel textures with the following SDS loss:
$\nabla_{{\psi}_{a}}\mathcal{L}^{{oc}}_{SDS}=\mathbb{E}_{t,\epsilon}\left[({\epsilon}_{{\phi}_{oc}}(c^{a}_{t};d^{a},y_{a},t)-\epsilon)\frac{\partial c^{a}}{\partial{\psi}_{a}}\right],$ (17)
$\nabla_{{\psi}_{a}}\mathcal{L}^{{sdc}}_{SDS}=\mathbb{E}_{t,\epsilon}\left[({\epsilon}_{{\phi}_{sd}}(c^{a}_{t};y_{a},t)-\epsilon)\frac{\partial c^{a}}{\partial{\psi}_{a}}\right],$ (18)
where $c^{a}$ is the color image. Additionally, human-specific diffusion models are combined to optimize the appearance of outfits in the dressed state to ensure basic texture harmony, and the SDS loss is formulated as:
$\nabla_{{\psi}_{a}}\mathcal{L}^{{hc}}_{SDS}=\mathbb{E}_{t,\epsilon}\left[({\epsilon}_{{\phi}_{hc}}(c^{a+h}_{t};n^{a+h},y_{a+h},t)-\epsilon)\frac{\partial c^{a+h}}{\partial{\psi}_{a}}\right].$ (19)
In summary, the loss function for optimizing the apparel appearance is as follows:
$\mathcal{L}_{{app-tex}}=\mathcal{L}^{{hc}}_{SDS}+\mathcal{L}^{{oc}}_{SDS}+\mathcal{L}^{{sdc}}_{SDS}.$ (20)
III-E Unified Texture Refinement
After the above steps, we produce detailed geometry and lifelike textures for human bodies, garments, shoes, and accessories. However, a domain gap in the training data for fine-tuning expert models causes texture disharmony between the body and outfits, reducing realism (Fig. 8-(c)). To address this issue, we propose a unified texture refinement (UTR) strategy. Specifically, we fine-tune all apparel texture fields ${\psi}_{A}=\{\psi_{{a}_{i}},i\in[1,...,N]\}$ of the assembled avatar under the human-specific normal-conditioned diffusion model and the MSDS loss from Eq.8. The UTR strategy ensures visual unity and improves texture harmony across the avatar. It reduces the domain gap, enhancing realism and aesthetic quality.
III-F Implementation Details
Our algorithm is implemented using PyTorch [50] and ThreeStudio [17]. Generating a human body takes $\sim$3 hours (24GB VRAM); generating an outfit takes $\sim$4 hours (40GB VRAM). We use MLP with multi-resolution hash encoding [47] to efficiently represent SDF, mSDF, and color fields. Given a text describing a dressed human, e.g., “A man wearing jeans and a T-shirt”, we first use a rule-based script to decompose the full input text into multiple sub-prompts: “A man in his underwear”, “A pair of jeans”, and “A T-shirt”. We then generate a base human body based on “A man in his underwear”. Next, we generate jeans based on “A pair of jeans” and “A man wearing jeans”, followed by generating the T-shirt based on “A T-shirt” and “A man wearing jeans and a T-shirt”. Finally, fine-tune the texture of the assembled avatar based on the full input text “A man wearing jeans and a T-shirt”. Additional details can be found in the Supp. Mat.
Figure 6: Qualitative comparisons with baseline text-to-avatar and text-to-apparel methods. GarmentDreamer [31] and SO-SMPL [66] only support clothing generation.
Figure 7: Qualitative comparisons with the text-to-disentangled-avatar work SO-SMPL [66]. SO-SMPL not only exhibits significantly lower generation quality than our method but also fails to produce the diverse outfits shown in Fig. 5, as achieved by our Barbie-style avatars.
IV Experiments
Some examples of 3D bodies and outfits generated by our method are shown in Fig. 5. Due to space limitations, only the key experimental settings and results are presented here. For complete results and details, please refer to our Supp. Mat.
IV-A Experimental Settings
Baselines. We perform both quantitative and qualitative comparisons of DreamBarbie with state-of-the-art (SOTA) SDS-based methods for text-driven avatar generation and apparel generation. For avatar generation, we compare with text-to-holistic-avatar methods (DreamWaltz [23], TADA [39], X-Oscar [45], HumanGaussian [41], HumanNorm [22], and DreamWaltz-G [24]) and text-to-decoupled-avatar method (SO-SMPL [66]). For text-to-apparel generation, we compare with text-to-3D methods (MVDream [62], GaussianDreamer [75], RichDreamer [54], LucidDreamer [38]), text-to-garment method (GarmentDreamer [31]), and SO-SMPL [66].
TABLE II: Quantitative comparisons with baseline methods
Method BLIP-VQA $\uparrow$ BLIP2-VQA $\uparrow$ GQP $\uparrow$ TAP $\uparrow$
DreamWaltz 0.5819 0.5333 2.80 2.31
TADA 0.5306 0.5333 3.25 2.91
X-Oscar 0.5069 0.5292 3.43 3.62
HumanGaussian 0.6222 0.5069 2.35 2.05
HumanNorm 0.5611 0.5292 3.81 3.51
DreamWaltz-G 0.6444 0.5722 2.87 2.54
SO-SMPL 0.6500 0.6000 0.45 0.71
DreamBarbie (Ours) 0.7167/0.7333 0.6333/0.7000 81.04 82.35
MVDream 0.7000 0.5933 9.03 8.92
GaussianDreamer 0.7017 0.5283 3.84 4.14
RichDreamer 0.8100 0.6667 3.25 2.95
LucidDreamer 0.7533 0.6400 4.59 3.77
GarmentDreamer 0.6500 0.4500 0.78 0.63
SO-SMPL 0.8500 0.7667 0.26 0.37
DreamBarbie (Ours) 0.8667/0.9667 0.7667/0.9167 78.25 79.22
  • The best result is highlighted in bolded. GQP: generation quality preference (%), TAP: text-image alignment preference (%). $\dagger$: When calculating VQA metrics, we ignore the problem of retrieving shoes and accessories, as GarmentDreamer [31] and SO-SMPL [66] only support clothing generation.
[…]
Figure 8: Ablation Study of DreamBarbie.
Evaluation Metrics. Existing text-to-3D approaches utilize CLIP-based metrics to evaluate text-image alignment and compare the quality of generated results. However, CLIP-based metrics have been shown [20, 44] to be insufficient to accurately measure the fine-grained correspondence between 3D content and input prompts, which is further confirmed by experiments in the Supp. Mat.
Consequently, inspired by Progressive3D [9], we adopt fine-grained text-to-image evaluation metrics, including BLIP-VQA [34, 32] and BLIP2-VQA [33, 32], to evaluate the generation capacity of current methods and DreamBarbie. Specifically, we first convert the prompt into multiple separate questions to retrieve corresponding content, then feed the rendered image of the generated content into the VQA model and ask questions one by one, and finally use the probability of answering “yes” as the evaluation metric. For instance, the input avatar prompt “A man wearing X1.” is converted into “Is the person in the picture a man?” and “Is the person in the picture wearing X1?”. The input apparel prompt “A pair of X1.” is converted into “Is the object in the picture a pair of X1?”. Besides, we randomly select 40 cases (20 for avatar and 20 for outfit) to conduct a user study and ask 67 volunteers to assess (1) generation quality and (2) text-image alignment, and select the preferred methods.
IV-B Comparisons of Avatar Generation
Comparison with Text-to-Avatar Methods. As shown in the upper part of Table II, our method significantly outperforms all baseline methods across all evaluation metrics. The qualitative results presented in the left part of Fig. 6 further validate its superiority. Compared to existing methods, DreamBarbie generates avatars with more detailed and realistic geometry, while achieving better alignment with the input text, without omitting any specified apparel items. Moreover, DreamBarbie achieves a high degree of disentanglement among body, garments, shoes, and accessories, enabling flexible outfit combinations and edits, much like dressing physical Barbie dolls (Fig. 9-(a) and -(b)). Furthermore, thanks to the use of G-Shell and the proposed SMPLX-evolving prior loss, DreamBarbie supports expressive animation and compatibility with physical simulation (Fig. 9-(c) and -(d)).
Comparison with Text-to-Decoupled-Avatar Method. As illustrated in Fig. 7, our method surpasses SO-SMPL [66] in generating finer geometric details and more lifelike textures. Additionally, DreamBarbie enables the generation of various accessories (e.g., necklaces, glasses, and watches), which are not supported by SO-SMPL [66].
IV-C Comparisons of Apparel Generation
As shown in the lower part of Table II, our approach consistently outperforms all competing methods on all evaluation metrics. Qualitative comparisons are shown in the right part of Fig. 6, displaying representative generated outfits. Compared to other methods, DreamBarbie produces apparel with higher-fidelity geometry and textures, while maintaining accurate alignment with the input text, without introducing irrelevant elements such as body parts. Thanks to the expressive G-Shell [42] representation, we are able to accurately model the non-watertight topology of garments, enabling compatibility with physical simulations (Fig. 9-(d)).
IV-D Ablation Study
We conduct detailed ablation studies in Fig. 8 to validate the effectiveness of each component in DreamBarbie.
Effect of SMPLX-Evolving Prior Loss. As shown in Fig. 8-(a), omitting $\mathcal{L}_{prior}$ leads to distorted body proportions and unnatural hands, significantly degrading realism. Using a fixed human prior ensures anatomical plausibility but results in overly smooth surfaces with limited detail (e.g., missing muscle contours and hair). In contrast, the full ${L}_{prior}$ enables both reasonable and detailed human body generation while providing rich semantics that support downstream tasks such as apparel transfer, editing, animation, and simulation.
Effect of Dual-Space Optimization. To generate fine-detailed, well-fitted apparel, we employ dual-space optimization with human- and object-specific generative priors. As shown in Fig. 8-(b), human-space-only optimization yields assets with poor detail (orange box) and erroneously absorbs body parts into clothing (blue box), breaking semantic boundaries. In contrast, apparel-space-only optimization produces ill-fitting outfits—evident in oversized shoes (green box) and texture discontinuities near the waist (purple box)—undermining realism. Our dual-space approach jointly leverages both priors to achieve detailed, body-conforming garments and visually coherent avatars.
Effect of Unified Texture Refinement. To reduce visual inconsistency between the body and outfits, we introduce a unified texture refinement (UTR) strategy. Fig. 8-(c) confirms that UTR effectively reduces texture conflicts across components, thereby enhancing overall realism and aesthetic coherence.
Effect of Template-Preserving Loss. To preserve structural integrity in the generated apparel, we incorporate a template-preserving loss $\mathcal{L}_{{temp}}$. As shown in Fig. 8-(d), this loss is essential for preventing unexpected holes and other geometric artifacts that severely degrade apparel quality.
Effect of Hole-Preserving Loss. To accurately model intended garment openings (e.g., necklines), we propose $\mathcal{L}_{{hole}}$. As demonstrated in Fig. 8-(e), $\mathcal{L}_{{hole}}$ avoids floating triangles within holes, resulting in clean neckline structures.
Effect of Collision Loss. To avoid intersections between apparel and the underlying body, we introduce $\mathcal{L}_{{colli}}$. As illustrated in Fig. 8-(f), neglecting $\mathcal{L}_{{colli}}$ causes severe collision artifacts, which notably compromise result realism.
Figure 9: Applications of DreamBarbie.
IV-E Applications
As shown in Fig. 9-(a), DreamBarbie’s fine-grained decoupling enables the flexible combination of any human body with any outfit, similar to dressing physical Barbie dolls. Thanks to the SMPLX-evolving prior loss $\mathcal{L}_{prior}$, we achieve accurate alignment between the generated assets and the SMPL-X model. This alignment allows for intuitive apparel editing by adjusting shape parameters, and supports natural full-body animation through control of pose and expression parameters, as shown in Fig. 9-(b) and -(c). The expression sequences and body motions are taken from AIST++ [35] and TalkShow [74]. Furthermore, leveraging the expressive G-Shell [42] representation, our method accurately captures non-watertight garment topology, supporting physical simulation [63] as shown in Fig. 9-(d). Our method enables the creation of Barbie-style 3D avatar meshes. As a result, users can view and combine human bodies and outfits within professional 3D software [25], as illustrated in Fig. 9-(e).
Figure 10: Failure cases.
V limitations and Future Work
Limited by the resolution of G-Shell, our method struggles to accurately model highly complex geometries such as hair, ears, and other facial features, as shown in Fig. 10. In addition, while the use of predefined templates ensures reasonable initialization, it may slightly limit generation diversity. In future work, we aim to explore template-free, disentangled generation by leveraging 3D generative priors [72]. Finally, our reliance on the SDS loss introduces significant computational overhead, limiting practical applicability. As a direction for future work, we plan to adopt a feed-forward, disentangled generation framework such as StdGen [18], integrating diverse human datasets and advanced priors to enable efficient 3D Barbie-style avatar generation.
VI Conclusion
We propose DreamBarbie, a novel text-guided framework for creating an animatable 3D avatar dressed in decoupled shoes and accessories, along with simulation-ready garments, resembling iconic Barbie dolls. Specifically, we employ an expressive G-Shell to uniformly represent both watertight and non-watertight meshes. We introduce an SDF-based initialization that reformulates boundary definition from manifold geodesics into 3D field intersections, significantly enhancing stability and efficiency of open-surface modeling. To guarantee the domain-specific fidelity of the generated human body and outfits, we suitably combine different specific expert T2I models for domain-specific knowledge. To mitigate the negative impacts caused by the over-strong generative priors of the expert models, we propose a series of solid regularization losses and optimization strategies to ensure geometric rationality and texture harmony of the generated results. Extensive experiments demonstrate that our approach not only outperforms compared methods in both dressed avatar and apparel generation tasks, but also facilitates seamless composition and editing of outfits, as well as expressive animation and physical simulation.